ANNOUNCEMENTS ------------- * It may be impossible to access MDI data Saturday, 05 January since they are re-arranging the computer room at Stanford. * DSS-16 is in maintenance and won't be up until week 2 or so of 2002. * LASCO and EIT running difference images in java can be seen at: http://lasco-www.nrl.navy.mil/javagif/ FOT REPORT ---------- Spacecraft Status: Nominal Spacecraft Anomalies: None Accomplished Activities: VIRGO, ACU/CDMU Memory Dumps Planned Activities: VIRGO, NOM HGA Table, New tracking star (7.1 mag), CEPAC ESU Reset Upcoming Operations: None Ground Anomalies: 002/2029 - TC software anomalies at D27 (2 min delay) 003/0105 - Wayward command at D46 (4 min delay) SOLAR STATUS ------------ 01/02 DOY 002 - 01/03 DOY 003 EIT daily On-Line Solar Status and observations: http://umbra.nascom.nasa.gov/eit/plan/log_form20020103.html LASCO daily On-Line Solar Status and observations: http://lasco-www.nrl.navy.mil/observations/2001/Todays_obs.html Web Page for Planning: http://www.bbso.njit.edu/arm INSTRUMENT STATUS ----------------- CDS: Nominal. Supporting JOP151 on the east limb and lubrication study at 20:30 for 2 hours today. Friday - JOP151 on east limb. UVCS: Nominal. JOP151 on AR 9733 as it returns on the east limb, followed by mini synoptics. Friday - After mini synoptics finish at 7 UT UVCS returns to JOP151 observations on the east limb until 17 UT, when 24-hour synoptics begin. LASCO: Nominal. C2/C3 synoptics supporting JOP151. For the next month, color sequences will be taken daily in C3 for flat fielding purposes. EIT: Nominal. 195 Angstrom CME watch, synoptics, and JOP151. MDI: Nominal. Full disk campaign of magnetograms and dopplergrams. Supporting JOP151. Monthly calibrations are planned for this afternoon. TRACE: Nominal. Observing AR 9767 today.